OpenStudy (anonymous):

How do I do, (z^8)^0 z^5 ?

OpenStudy (anonymous):

Jack_RM i dont know sorry

OpenStudy (anonymous):

Anything to the power of 0 is 1. So (z^8)^0 = 1. 1*z^5 = z^5